# Guiding an Automatic Speech Recognition Decoder using Large Language Models

### v1

@@ VII-D Beam size and number of LLM candidates
[…]
@@ VII-E Analyzing acronyms
TABLE III: Examples of acronym transcriptions by different models. (V), (X), and ’–’ indicate correct, incorrect, and missing predictions, respectively.
Reference Greedy Beam 4-gram GCNN Transformer
NASA scheduled the launch of the space
shuttle discovery for september twenty ninth. NASA (V) scheduled A (X) scheduled NASA (V) scheduled - (X) scheduled
MICC said it intends to pay the dividend arrears
on july thirty first to stock of record july second. MICC (V) said I (X) said MICC (V) said I (X) said
RLI corporation a peoria illinois based insurance
holding company will begin trading friday on the
big board under the symbol RLI.
symbol RLI (V) corporation
symbol RLI (V)
- (X) corporation
symbol - (X)
@@
symbol - (X)
His MBA also irks some colleagues who
are contemptuous of foreign concepts. His MBA (V) also His - (X) also His MB (X) also His - (X) also
Two years ago BASF made three separate
acquisitions in the US.
@@
Model With Acronyms Without Acronyms
@@
Acronyms, words formed from the initial letters of a phrase and pronounced as individual characters (e.g., “U-S-A”), present a unique challenge for ASR systems. In WSJ0, acronyms are formatted with dots between letters (e.g., “u. s. a.”), though these dots are not valid symbols for the acoustic models’ character set. Although acoustic models can represent the phonetic sequences of acronyms, traditional n-gram language models often struggle with them. Not only do these models rely heavily on predefined vocabularies with limited ability to handle out-of-vocabulary (OOV) words, but they also output acronyms as single lowercase words, making it impossible to distinguish them from regular words in the text.
@@
This advantage is evident in the results shown in Table IV. Our method achieves a WER of 4.78% and CER of 0.98% on sentences with acronyms, outperforming all other baselines. As shown in Table V, our method achieves an acronym recognition accuracy of 0.88, outperforming the 4-gram (0.39), GCNN (0.70), and Transformer (0.27) baselines, and approaching the accuracy of the greedy decoder (0.93). Representative transcription examples illustrating these trends are provided in Table III. Representative examples of acronym transcriptions across decoding strategies are shown in Table III. =
@@ VIII Conclusions
In this work, we introduced a novel zero-shot decoding approach that positions LLMs as the primary driver of the decoding process in automatic ASR systems, moving away from traditional approaches dominated by acoustic models. We derived this approach from the MAP estimator of tokens given the speech signal and proposed an iterative procedure to solve it efficiently.
@@
Experiments across multiple ASR benchmarks demonstrated consistent improvements over strong baselines, validating the robustness and adaptability of the proposed approach. At a broader level, our goal is to close the performance gap between modular ASR systems with separately trained acoustic and language models and end-to-end supervised transformer-based models such as Whisper [34].
The proposed approach presents two significant drawbacks, which we will address in future research. Firstly, while LLMs demonstrate substantial advantages over traditional language models across various tasks, this does not correspond to a marked improvement in the WER of the proposed ASR system. We believe this issue stems from a relatively weak acoustic model. Future work will focus on integrating more robust acoustic models, such as the encoder from Whisper [34]. This integration should enhance both acoustic and language representations and provide a means to replace the internal, relatively weak language model of Whisper.
The second direction for future work involves adapting the current method for streaming ASR, which is crucial for real-time applications. This requires rethinking the system’s computational flow, particularly by modifying the acoustic model to support incremental processing. One promising avenue is to restructure the attention mechanism—typically a bottleneck in non-streaming models—so that attention matrices can be computed and updated incrementally, frame by frame. Such advancements would enable the ASR system to produce partial transcriptions on-the-fly while maintaining the benefits of deep integration with a powerful LLM.

### v2

@@ VII-D Beam size and number of LLM candidates
We evaluated the influence of the LLM beam size and the number of token candidates from the LLMs. The optimal configuration parameters were determined through empirical analysis, as shown in Figures 2 and 2. A beam size of 5 and top-$k$ value of 5,000 were selected based on the performance plateaus observed in these analyses. The dataset-specific hyperparameters ($\alpha$, $\beta$) were tuned for each language model-dataset combination, as detailed in Table I, maintaining consistency across acoustic models due to their shared CTC training objective.
@@ VII-E Analyzing acronyms
TABLE III: Examples of acronym transcriptions by different models. (V), (X), and ’-’ indicate correct, incorrect, and missing predictions, respectively.
Reference Greedy Beam 4-gram GCNN Transformer Ours
NASA scheduled the launch of the space shuttle discovery for september twenty ninth. NASA (V) scheduled A (X) scheduled NASA (V) scheduled (X) scheduled NASA (V) scheduled
MICC said it intends to pay the dividend arrears on july thirty first to stock of record july second. MICC (V) said I (X) said MICC (V) said I (X) said MICC (V) said
RLI corporation a peoria illinois based insurance holding company will begin trading friday on the big board under the symbol RLI.
@@
symbol (X)
RLI (V) corporation
symbol RLI (V)
(X) corporation
symbol (X)
RLI (V) corporation
symbol RLI (V)
His MBA also irks some colleagues who are contemptuous of foreign concepts. His MBA (V) also His (X) also His MB (X) also His (X) also His MBA (V) also
Two years ago BASF made three separate acquisitions in the US.
@@
Model w/ Acronyms w/o Acronyms
@@
Acronyms, words formed from the initial letters of a phrase and pronounced as individual characters (e.g., ”U-S-A”), present a unique challenge for ASR systems. In WSJ0, acronyms are formatted with dots between letters (e.g., ”u. s. a.”), though these dots are not valid symbols for the acoustic models’ character set. Although acoustic models can represent the phonetic sequences of acronyms, traditional n-gram language models often struggle with them. Not only do these models rely heavily on predefined vocabularies with limited ability to handle out-of-vocabulary (OOV) words, but they also output acronyms as single lowercase words, making it impossible to distinguish them from regular words in the text.
@@
This advantage is evident in the results shown in Table IV. Our method achieves a WER of 4.78 and CER of 0.98 on sentences with acronyms, outperforming all other baselines. As shown in Table V, our method achieves an acronym recognition accuracy of 0.88, outperforming the 4-gram (0.39), GCNN (0.70), and Transformer (0.27) baselines, and approaching the accuracy of the greedy decoder (0.93). Representative transcription examples illustrating these trends are provided in Table III. Representative examples of acronym transcriptions across decoding strategies are shown in Table III.
VII-F Semantic fidelity evaluation
To complement the word-level accuracy results, we evaluated the semantic fidelity of each decoding strategy using the BERTScore $F_{\text{BERT}}$ measure [zhang2019bertscore], which computes token-level semantic similarity based on contextualized BERT embeddings. Table VI reports results on WSJ0, TEDLIUM, and ALLSTAR-ENG, with WSJ0 further split into sentences with and without acronyms. The proposed decoder achieves the highest $F_{\text{BERT}}$ across all datasets, with the largest gains on acronym-containing sentences where conventional decoders often mis-segment or expand abbreviations.
Wilcoxon signed-rank tests confirm that these improvements are statistically significant (WSJ0: $p<0.001$; TEDLIUM: $p<0.001$; ALLSTAR-ENG: $p<0.0001$). Even under matched-WER conditions ($|\Delta\text{WER}|\leq 0.01$), the proposed decoder maintains positive semantic gains (e.g., WSJ0 acronym subset: $\Delta F_{\text{BERT}}=0.008$-$0.014$, $p<0.01$), indicating improved meaning preservation beyond what is captured by word-level metrics alone.
TABLE VI: Semantic fidelity comparison using BERTScore $F_{\text{BERT}}$ across WSJ0, TEDLIUM, and ALLSTAR-ENG. For WSJ0, results are additionally reported for sentences with and without acronyms. The proposed decoder achieves the highest $F_{\text{BERT}}$ on all datasets, reflecting improved semantic similarity to the reference transcripts.
WSJ0 TEDLIUM ALLSTAR-ENG
Model w/ Acronyms w/o Acronyms Overall
Greedy 0.9778 0.9870 0.9862 0.9764 0.9905
4-gram 0.9645 0.9874 0.9852 0.9767 0.9918
GCNN 0.9755 0.9866 0.9856 0.9831 0.9935
Transformer 0.9620 0.9880 0.9863 0.9832 0.9936
Ours 0.9861 0.9897 0.9892 0.9850 0.9949
VII-G Error-correction baselines
The results above focus on integrating language models directly into the decoding process. A natural question is how such decoding-time integration compares with LLM-based post-hoc error-correction approaches that operate on completed ASR hypotheses. To examine this, we evaluate several zero-shot error-correction baselines applied to the outputs of the HuBERT X-Large ASR system and compare them with our LLM-guided decoder.
For consistency and reproducibility, all post-hoc error-correction baselines are implemented using open-weight Llama-3 models [dubey2024llama]. The instruction-tuned variant is used for hypothesis generation, and the base model is used for scoring where applicable, following the same zero-shot prompting baseline setups.
We implement the zero-shot error-correction baselines proposed by Ma et al. [ma2023can]: a zero-shot, unconstrained setting, in which the LLM receives an ASR $N$-best list and generates a corrected transcription; a zero-shot selective rescoring setting, where the LLM is instructed to select a single hypothesis using a constrained XML-style interface; and a one-shot, unconstrained variant that augments the prompt with a single in-context example. Additionally, we implement the post-hoc baseline N-best Closest Decoding [ma2024asr], where the unconstrained LLM-generated correction is mapped to the ASR hypothesis with the minimum Levenshtein distance. We also include a standard $N$-best rescoring baseline that reorders ASR hypotheses by LLM scores and selects the top-ranked hypothesis.
In addition, we evaluate ProGRes [tur2024progres], which uses a multi-step post-hoc pipeline with two different LLMs applied to the ASR $N$-best list. First, the acoustic model produces an $N$-best list using CTC beam search. An instruction-tuned Llama-3 model then generates an additional hypothesis conditioned on this list. Finally, ProGRes computes language pseudo-log-likelihoods for each hypothesis using a separate Llama-3 base model, recomputes post-hoc acoustic scores via an additional CTC log-likelihood pass, and selects the final output through linear fusion of these scores. Following their setup, the LLM score interpolation weight $\alpha$ is tuned on development data. The optimal value is $\alpha=0.1$ for ALLSSTAR-Eng, whereas for WSJ0 and TED-LIUM 3 the best performance is obtained with $\alpha=0$, meaning the LLM-based score does not affect the selected hypothesis for these datasets.
Overall, results are summarized in Table VII. The table reports WER and CER, along with word insertion rates (Ins%) to reflect each method’s tendency to introduce content not supported by the acoustic evidence. We additionally report relative inference overhead, which reflects the number of extra inference operations beyond the common CTC beam-search decoding step, expressed in terms of LLM hypothesis generation, LLM scoring over an ASR $N$-best list, and additional CTC forward passes. For our decoder, integrated scoring amortizes this overhead via prefix reuse and KV caching (Section VII-B).
While some post-hoc baselines achieve competitive performance on ALLSSTAR-Eng, which consists of short, read utterances, performance differences become more pronounced on WSJ0 and TED-LIUM 3, which represent more complex domains with longer, structured text and spontaneous speech, respectively. Post-hoc methods do not condition generation on the acoustic signal and may therefore produce semantically reasonable outputs that are not supported by the audio, commonly referred to as hallucinations. For example, the selective rescoring variant proposed by Ma et al. performs poorly, with extremely high WER and CER, because instruction-tuned LLMs do not always follow output-format constraints. This behavior is also observed for ProGRes, which, despite achieving the lowest WER and CER on the ALLSSTAR-Eng dataset, produces hallucinations such as extending ”is the kangerlusig river in southwest” to ”is the kangerlusig river in southwest Greenland,” or augmenting ”they waited for an hour” with templated continuations that append spurious metadata. This trend is also reflected in higher insertion rates for post-hoc methods, as reported in Table VII, indicating a greater tendency to introduce acoustically unsupported content.
TABLE VII: Error-correction performance across ALLSSTAR-Eng, WSJ0, and TED-LIUM 3. WER, CER, word insertion rate (Ins%), and relative inference overhead are reported. All methods start from ASR CTC beam-search decoding.
Method
Relative inference
overhead
ALLSSTAR Eng WSJ0 TED-LIUM 3
WER (Ins%) $\downarrow$ CER $\downarrow$ WER (Ins%) $\downarrow$ CER $\downarrow$ WER (Ins%) $\downarrow$ CER $\downarrow$
[…]
ProGRes [tur2024progres]
[…]
$N\times$ LLM scoring +
$N\!\times\!\!$ CTC pass
3.24 (0.31) 1.92 8.46 (1.46) 1.65 9.86 (1.90) 3.32
N-best rescoring [chelba2012large] $N\times$ LLM scoring 4.29 (0.49) 2.26 6.72 (0.61) 3.47 14.91 (1.24) 11.01
Ours
integrated LLM scoring
with KV caching
during beam search
4.67 (0.68) 2.26 4.04 (1.01) 0.72 6.81 (1.62) 2.73
To quantify this systematically, we use the hallucination identification setup of Frieske and Shi [frieske2024hallucinations], which flags utterances with WER above 30%, cosine similarity below 0.2, and language-model perplexity below 200. The resulting sentence-level hallucination rates are reported in Table VIII. Our decoder maintains near-zero hallucination rates across all datasets, whereas post-hoc methods may hallucinate, with the largest effect observed for the selective interface baseline. Perplexity is computed using a Flan-T5-small language model, and semantic similarity is measured using sentence embeddings from all-MiniLM-L6-v2
TABLE VIII: Sentence-level hallucination rates across ALLSSTAR-Eng, WSJ0, and TED-LIUM 3. Hall.% (N) reports the percentage and absolute count of hallucinated utterances.
Method ALLSSTAR Eng WSJ0 TED-LIUM 3
Hall.% (N) $\downarrow$ Hall.% (N) $\downarrow$ Hall.% (N) $\downarrow$
0-shot uncon [ma2023can] 0.00 (0) 0.00 (0) 0.35 (4)
0-shot select [ma2023can] 13.19 (364) 13.57 (90) 22.60 (261)
1-shot uncon [ma2023can] 0.00 (0) 0.00 (0) 0.35 (4)
N-best Closest Decoding [ma2024asr] 0.00 (0) 0.00 (0) 0.35 (4)
N-best rescoring [chelba2012large] 0.40 (11) 2.56 (17) 6.58 (76)
ProGRes [tur2024progres] 0.40 (11) 0.00 (0) 0.35 (4)
Ours 0.00 (0) 0.00 (0) 0.35 (4)
@@ VIII Conclusions
In this work, we introduced a novel zero-shot decoding approach that positions LLMs as the primary driver of the decoding process in SSL-ASR systems, moving away from traditional approaches dominated by acoustic models. We derived this approach theoretically from the MAP estimator of tokens given the speech signal and proposed an iterative procedure to solve it efficiently.
@@
In contrast to post-hoc error correction and N-best, rescoring approaches, which apply an LLM only to the final ASR output, our method integrates the LLM directly into the decoding process. This enables a step-by-step interaction between linguistic modeling and acoustic evidence, favoring hypotheses that remain aligned with the speech signal and avoiding acoustically unsupported additions.
Experiments across multiple benchmarks demonstrated improved semantic fidelity and competitive WER relative to strong baselines, validating the robustness and adaptability of the proposed approach. At a broader level, our goal is to close the performance gap between modular ASR systems with separately trained acoustic and language models and end-to-end supervised transformer-based models such as Whisper [radford2023whisper].
The proposed approach presents two significant drawbacks, which we will address in future research. Firstly, while LLMs demonstrate substantial advantages over traditional language models across various tasks, this does not correspond to a marked improvement in the WER of the proposed ASR system. We believe this issue stems from a relatively weak acoustic model. Future work will focus on integrating more robust acoustic models, such as the encoder from Whisper [radford2023whisper]. This integration should enhance both acoustic and language representations and provide a means to replace the internal, relatively weak language model of Whisper.
The second direction for future work involves adapting the current method for streaming ASR, which is crucial for real-time applications. This requires rethinking the system’s computational flow, particularly by modifying the acoustic model to support incremental processing. One promising avenue is to restructure the attention mechanism-typically a bottleneck in non-streaming models-so that attention matrices can be computed and updated incrementally, frame by frame. Such advancements would enable the ASR system to produce partial transcriptions on-the-fly while maintaining the benefits of deep integration with a powerful LLM.